# Quality or Quantity? On Data Scale and Diversity in Adapting Large Language Models for Low-Resource Translation

### v1

@@ Abstract
Despite the recent popularity of Large Language Models (LLMs) in Machine Translation (MT), their performance in low-resource translation still lags significantly behind Neural Machine Translation (NMT) models. In this paper, we explore what it would take to adapt LLMs for low-resource settings. In particular, we re-examine the role of two factors: a) the importance and application of parallel data, and b) diversity in Supervised Fine-Tuning (SFT). Recently, parallel data has been shown to be less important for MT using LLMs than in previous MT research. Similarly, diversity during SFT has been shown to promote significant transfer in LLMs across languages and tasks. However, for low-resource LLM-MT, we show that the opposite is true for both of these considerations: a) parallel data is critical during both pretraining and SFT, and b) diversity tends to cause interference, not transfer. Our experiments, conducted with 3 LLMs across 2 low-resourced language groups - indigenous American and North-East Indian - reveal consistent patterns in both cases, underscoring the generalizability of our findings. We believe these insights will be valuable for scaling to massively multilingual LLM-MT models that can effectively serve lower-resource languages.
@@ 1 Introduction
Large Language Models (LLMs) have been dominating recent research in Machine Translation (MT), showing good few-shot prompting (Garcia et al., 2023; Hendy et al., 2023) and stronger instruction-tuning (Alves et al., 2024; Xu et al., 2024) performances – even outperforming conventional Neural Machine Translation (NMT) models, trained mainly on parallel data, in the latest WMT shared task (Kocmi et al., 2024). However, LLM translation for low-resource languages (LRLs) still lags significantly behind NMT models (Robinson et al., 2023; Zhu et al., 2024b). While the strong performance of LLMs on high-resource languages can be attributed to the skewed language distribution during pre-training and the unintentional consumption of parallel data at scale (Briakou et al., 2023), no such relief exists for LRLs. This leads to the main question motivating this paper: What would it take to adapt LLMs for low-resource MT?
@@
Table 1: Comparing data scales of previous works with ours, in terms of pretraining token counts, the base LLM (if pretraining continued from one) and how many languages this spanned. Estimated from the reported sentence count assuming 100 tokens per sentence.
Recently, the role of parallel data at scale, long viewed as fundamental to the success of NMT models, has come into question in the era of LLM-MT systems. Motivated by the modest gains of training on 300M parallel sentences (Yang et al., 2023), and the surprising benefits of scaling down during SFT (Zhou et al., 2023), subsequent works have used only tens of thousands of human-written bitext for LLM-MT (Zhang et al., 2023b; Alves et al., 2024; Xu et al., 2024) – with SFT scaling laws further showing the early plateau of LLM-MT performance (Zhang et al., 2024a). Even more surprisingly, Zhu et al. (2024a) showed MT abilities emerging with just 32 SFT examples! However, these explorations concern LLMs pre-trained on several billions of tokens in the languages in question. We revisit these notions for low-resource MT and work with languages having datasets that are 2-3 orders of magnitude smallerTable 1 compares the scale of the datasets involved in our work and related research. We discover that for low-resource MT, parallel data is critical not just during CPT, but even more so during SFT - in direct contrast with research on high-resource languages (HRLs).
Next, diversity in tasks, prompts, and datasets during SFT has been shown to significantly improve model performance across a range of tasks (Mishra et al., 2022; Chung et al., 2024). MT instructions have been shown to majorly improve diverse multilingual tasks (Ranaldi and Pucci, 2023; Zhu et al., 2023; Alves et al., 2024). With regard to linguistic diversity, Muennighoff et al. (2023) showed that MT data in seen languages can help boost unseen language performance. Inspired by these findings, we study if diversity could truly enrich SFT if our goal is solely low-resource MT for a fixed set of pairs. By conducting experiments across a range of tasks and MT directions with SFT datasets of varying compositions, we establish that diversity leads to negative interference, and fine-tuning on multilingual MT is the optimal strategy. In fact, we observe that training for more epochs on MT data is more effective than curating and training on a diverse dataset of the same size.
@@
To ensure the generalizability of our findings, we conduct 2 sets of experiments training multilingual LLMs on different sets of languages: a) 11 indigenous American and b) 4 North East Indian languages, wherein the former follows a Latin script and the latter has some languages that do not. Our focus in this work is on the eng/spa $\rightarrow$ X direction, since generation in an LRL is known to be a much harder task than in an HRL like English or Spanish, and we are interested in studying the challenges involved. We experiment with 3 base LLMs of varying sizes – Gemma 2B (Gemma Team et al., 2024), Mistral 7B (Jiang et al., 2023), and Llama 3 8B (Dubey et al., 2024), and observe that findings are mostly consistent across these models.
By applying our findings to 2-stage training, our methods achieved a +16.5 average chrF++ improvement over few-shot prompting – with the largest gains coming from the 8 least-resourced American languages in our setup, all of which have about 10K-50K parallel sentences each. We hope that the findings of this work would be useful when scaling to LLMs that can effectively translate into lower-resource languages.
@@ 2 Related Work
High-Resource Translation LLMs
There has been considerable interest in using LLMs as MT systems recently. Following initial success in prompting LLMs for high-resourced pairs (Vilar et al., 2023; Garcia et al., 2023; Hendy et al., 2023; Zhang et al., 2023a; Iyer et al., 2023), subsequent works have attempted to train LLMs on parallel data at scale (Yang et al., 2023; Lu et al., 2024), but these yielded modest gains and underperformed smaller encoder-decoder baselines such as NLLB-200 (Costa-jussà et al., 2022). Zhang et al. (2024a) showed through scaling laws for SFT that LLMs pre-trained at the order of 50B-300B tokens saturate in MT performance with 20K-30K instructions. Concurrently, (Xu et al., 2024) discovered excess parallel data washed out LLM knowledge, so they proposed a 2-stage paradigm called ALMA that involved pre-training on scaled-up monolingual data, followed by SFT on much smaller high-quality bitext (~60K lines). ALMA outperformed NLLB-200. Following their success, Alves et al. (2024) adopted the ALMA framework to train Tower 7B for 10 high-resourced languages outperformed ALMA and also matched GPT-4.
It is worth noting that Xu et al. (2024) did not include parallel data during CPT. However, inspired by research showing LLMs unintentionally consume parallel data at scale (Briakou et al., 2023), several works have included it at the order of several billions of tokens for the top 10-20 high-resource languages (Anil et al., 2023; Wei et al., 2023; Fujii et al., 2024; Alves et al., 2024). Concurrent work has explored pre-training on synthetic, code-switched parallel data for 101 languages (Lu et al., 2024) with a total of 400M sentences This work explores the impact of parallel data exclusively for low-resource performance and experiments with 1M–13M parallel sentences (50M–750M tokens) during pre-training, two orders of magnitude smaller than prior work. Moreover, we discover that bitext is particularly important for the least-resourced languages in our setup, which enjoys the highest and most consistent gains.
@@
Evaluation in massively multilingual settings have shown that LLMs are poor at low-resource translation (Robinson et al., 2023; Zhu et al., 2024b). In response, some works have advocated leveraging external resources in the MT pipeline, including multilingual lexicons (Lu et al., 2023), rule-based linguistic tools (Zhang et al., 2024b), word alignments (Mao and Yu, 2024) and even entire grammar books (Reid et al., 2024). However, such approaches create dependencies on resources and hinder extensibility across languages. Instead, we focus on extensibility and examine effective ways of utilising available data through CPT and SFT.
@@
There has been a body of work exploring multilingual instruction tuning that touched on diversity and data scale, but most of it was limited to HRLs. MT has been shown to improve cross-lingual generation (Ranaldi and Pucci, 2023; Zhu et al., 2023), while Chen et al. (2024) showed multilingual instruction tuning on machine-translated Alpaca datasets matches or beats monolingual tuning. Kew et al. (2023) and Shaham et al. (2024) showed a small quantity of multilingual instruction tuning data can improve cross-lingual generation capabilities in medium and high-resource languages, while Zhu et al. (2024a) showed only 32 examples in HRLs suffice to elicit MT capabilities from LLMs. In our work, we conclude that these findings have the common denominator of pre-training on HRLs at scale. When moving to LRLs, transfer is not as trivial as one might expect, and scaling task-specific data, without diversity, is the most effective solution.
@@ 3 Approach
We describe our efforts to adapt the widely-used ALMA framework – originally designed for fine-tuning LLMs for HRL translation (Xu et al., 2024; Alves et al., 2024) – for LRL translation.
@@ 3.1 Stage 1: Continued Pre-training (CPT)
In a scenario where monolingual data is scarce, it is crucial to investigate the most effective way to use parallel data – which, for our indigenous American languages, was found to surprisingly be more abundant than the former [^1] While monolingual online data is scarce, many translations of constitutions, articles etc. from Spanish do exist. We investigate 3 methods of mixing all available parallel and monolingual data:
@@
    All Mono: Here, we use monolingual data and the target side of all available parallel data, essentially using it as a source of additional monolingual data
@@
    Mono + parallel (concat): This too includes monolingual data, but also incorporates parallel data in the form of concatenated source-target pairs. We prepend source and target language codes respectively before concatenation and, following Guo et al. (2024), use a newline delimiter for separating them.
@@
    Mono + parallel (separate): This includes all monolingual data, and the source side and target side of parallel data as separate sentences, as opposed to concatenating them in one.
We depict our approach in Fig 1. In the first option, the motivation is that it might be hard for the LLM to learn to model concatenated sequences, given that they were likely scarce in the original pre-training corpus, with the added challenge of pre-training on ‘new’ low-resource languages. On the other hand, if the model is able to adapt to concatenated sequences, it could make the LLM more adjusted to the task of translation. Finally, the third method verifies whether the results of the ’concatenated’ model are due to the concatenation itself, rather than simply being exposed to additional tokens in the source language, as compared to the first option. We use this terminology throughout the experiments in our paper, but later shorten our best performing baseline ‘Mono + parallel (concat)’ to ‘Mono + parallel’ for brevity.
@@
While these techniques specify how parallel data is incorporated in pre-training, we are also interested in controlling the ratio of parallel to monolingual data in the corpus, particularly for the ‘concat’ method - given pre-training on 100% parallel data can be suboptimal (Alves et al., 2024). To get a desired mixing ratio for ‘concat’, we include bitext only until it comprises a given percentage of the training corpus. Once this threshold is crossed, we use the ‘all mono’ method to include it as monolingual data instead. Next, we also use temperature sampling (Arivazhagan et al., 2019) to control the language-wise distribution in our monolingual and parallel pre-training corpora, since these are quite heterogeneous – with certain languages being extremely low-resourced. We sample monolingual and parallel data independently if using the ‘concat’ method, else we just mix them all together and then shuffle it to create our final pre-training corpus.
@@ 3.2 Stage 2: Supervised Fine-Tuning (SFT)
Next, we fine-tune LoRA adapters on supervised instruction data. We detail the tasks explored during SFT below. Note that we convert all instructions to the standard Alpaca format, and compute loss on the target tokens only (Taori et al., 2023).
@@
Apart from MT data in the LRLs, we also experiment with adding HRL MT data since it is more abundant and of higher quality, known to be important during SFT (Xu et al., 2024). To explore the impact of transfer learning, we work with HRL data that is in some way related to the source/target language, for eg. Spanish-English data for experiments on Spanish-X. Instructions are formatted in the same way as the LRL data.
@@
4 Experiments
We fine-tune 2 separate sets of multilingual LLMs for 2 different language groups, so as to facilitate evaluation on test sets from 2 different low-resource MT shared tasks – AmericasNLP 2024 (Ebrahimi et al., 2024) and the Indic track of WMT 2023 (Pal et al., 2023). The former involves 11 Indigenous Central & South American, while the latter focuses on 4 North-East (NE) Indian languages. The first group includes Aymara (aym), Bribri (bzd), Asháninka (cni), Chatino (ctp), Guarani (grn), Huichol (hch), Nahuatl (nhe), Otomi (ote), Quechua (quy), Shipibo-Konibo (shp) and Tarahumara (tar). The second consists of Khasi (kha), Meitei (mni), Mizo (lus) and Assamese (asm). Our motivation in choosing these languages was to experiment with LRLs containing both Latin (American) and non-Latin (Indic) scripts; that also had widely used, high-quality test sets. We use the former for our main experiments, and replicate the most interesting baselines in the Indic languages.
@@ 4.1 Data
Table 3: Parallel data sentence counts, from source to each target language. HRL and LRL indicate the total sizes for the 3 high-resource languages (Aymara, Guarani and Quechua) and the other 8 low-resource languages respectively.
@@
We curate parallel data from various sources, for use in both CPT and SFT. Tables 3 and 4 show the sizes for the American and Indian languages respectively. Given that our primary exploration is for the American languages with limited spa-X data, we also sample eng-X and por-X from OPUS (Tiedemann, 2012). Note the heavily skewed language distribution, with the 3 HRLs constituting  80% of spa-X data and 96% of the overall data. The lesser skew for spa-X is due to the efforts of AmericasNLP to collect data for these pairs and the prevalence of Spanish in Latin American countries. A similar skew also exists for the Indic languages, with English-Khasi being the least-resourced pair. We list all sources for the collected data in Appendix A.

### v2

@@ Abstract
Despite the recent popularity of Large Language Models (LLMs) in Machine Translation (MT), their performance in low-resource languages (LRLs) still lags significantly behind Neural Machine Translation (NMT) models. In this work, we explore what it would take to adapt LLMs for the low-resource setting. Particularly, we re-examine the role of two factors: a) the importance and application of parallel data, and b) diversity in Supervised Fine-Tuning (SFT). Recently, parallel data has seen reduced use in adapting LLMs for MT, while data diversity has been embraced to promote transfer across languages and tasks. However, for low-resource LLM-MT, we show that the opposite is true for both considerations: a) parallel data is critical during both pre-training and SFT; b) diversity tends to cause interference instead of transfer. Our experiments with three LLMs across two low-resourced language groups—Indigenous American and North-East Indian—reveal consistent trends, underscoring the generalizability of our findings. We believe these insights will be valuable for scaling to massively multilingual LLM-MT models that can effectively serve LRLs.
@@ 1 Introduction
Large Language Models (LLMs) have been dominating recent research in Machine Translation (MT), showing good few-shot prompting (Garcia et al., 2023; Hendy et al., 2023) and stronger instruction-tuning (Alves et al., 2024; Xu et al., 2024) performances—recently even outperforming commercial Neural Machine Translation (NMT) models (Kocmi et al., 2024). However, LLM translation for low-resource languages (LRLs) still lags significantly behind NMT models (Robinson et al., 2023; Zhu et al., 2024b). While the strong performance of LLMs on high-resource languages can be attributed to the skewed language distribution during pre-training and the unintentional consumption of parallel data at scale (Briakou et al., 2023), no such relief exists for LRLs. This leads to the main question motivating this paper: What would it take to adapt LLMs for low-resource MT?
@@
Table 1: Comparing data scales of previous works with ours, in terms of pre-training token counts, the base LLM (if pre-training continued from one) and how many languages this spanned. Estimated from the reported sentence count assuming 100 tokens per sentence.
Recently, the role of parallel data at scale, long viewed as fundamental to the success of NMT models, has come into question in the era of LLM-MT systems. Motivated by the modest gains of training on 300M parallel sentences (Yang et al., 2023), and the surprising benefits of scaling down during SFT (Zhou et al., 2023), subsequent works have used only tens of thousands of human-written bitext for LLM-MT (Zhang et al., 2023b; Alves et al., 2024; Xu et al., 2024)—with SFT scaling laws further showing the early plateau of LLM-MT performance (Zhang et al., 2024a). Even more surprisingly, Zhu et al. (2024a) showed MT abilities emerging with just 32 SFT examples! However, these explorations concern LLMs pre-trained on several billions of tokens in the languages in question. We revisit these notions for low-resource MT and work with languages having datasets that are 2-3 orders of magnitude smaller. In Table 1 we compare the scale of the datasets used in our work and related research. We discover that for low-resource MT, parallel data is critical not just during CPT, but even more so during SFT—in direct contrast with research on high-resource languages (HRLs).
Next, diversity in tasks, prompts, and datasets during SFT has been shown to significantly improve model performance across a range of tasks (Mishra et al., 2022; Chung et al., 2024). MT instructions have been shown to not just boost translation performance in unseen languages (Muennighoff et al., 2023), but also enhance LLM capabilities across diverse multilingual generation tasks (Ranaldi and Pucci, 2023; Zhu et al., 2023). Inspired by these findings, we study if SFT diversity could benefit low-resource LLM-MT systems too. By conducting experiments across a range of tasks and language pairs with SFT datasets of varying compositions, we establish that diversity leads to negative interference and fine-tuning on multilingual MT is the optimal strategy. Further, we observe that training for more epochs on MT data is more effective than curating and training on a diverse dataset of the same size.
@@
To ensure the generalizability of our findings, we conduct 2 sets of experiments training multilingual LLMs on different sets of languages: a) 11 Indigenous American and b) 4 North East Indian languages, wherein the former follows a Latin script and the latter includes languages that do not. Our focus in this work is on the eng/spa $\rightarrow$ X directions since generation in an LRL is known to be a much harder task than in an HRL like English or Spanish, and we are interested in studying the challenges involved. We experiment with 3 base LLMs of varying sizes—Gemma 2B (Gemma Team et al., 2024), Mistral 7B (Jiang et al., 2023), and Llama 3 8B (Dubey et al., 2024), and observe that findings are mostly consistent across these models.
By applying our findings to 2-stage training, our methods achieved a +16.5 average chrF++ improvement over few-shot prompting—with the largest gains coming from the 8 least-resourced American languages in our setup, all of which have about 10K-50K parallel sentences each. We hope that the findings of this work will be useful when scaling to LLMs that can effectively translate into lower-resource languages.
@@ 2 Related Work
High-Resource Translation with LLMs
There has been considerable interest in using LLMs as MT systems recently. Following initial success in prompting LLMs for high-resourced pairs (Vilar et al., 2023; Garcia et al., 2023; Hendy et al., 2023; Zhang et al., 2023a; Iyer et al., 2023), subsequent works have attempted to train LLMs on parallel data at scale (Yang et al., 2023; Lu et al., 2024), but these yielded modest gains and underperformed smaller encoder-decoder baselines such as NLLB-200 (Costa-jussà et al., 2022). Zhang et al. (2024a) showed through scaling laws for SFT that LLMs pre-trained at the order of 50B-300B tokens saturate in MT performance with 20K-30K instructions. Concurrently, Xu et al. (2024) discovered excess parallel data washed out LLM knowledge, so they proposed a 2-stage paradigm called ALMA that involved pre-training on scaled-up monolingual data, followed by SFT on much smaller high-quality bitext (~60K lines). ALMA outperformed NLLB-200. Following their success, Alves et al. (2024) adopted the ALMA framework to train Tower 7B for 10 high-resourced languages, outperforming ALMA and also matching GPT-4.
It is worth noting that Xu et al. (2024) did not include parallel data during CPT. However, inspired by research showing LLMs unintentionally consume parallel data at scale (Briakou et al., 2023), several works have included it at the order of several billions of tokens for the top 10-20 high-resource languages (Anil et al., 2023; Wei et al., 2023; Fujii et al., 2024; Alves et al., 2024). Concurrent work has explored pre-training on synthetic, code-switched parallel data for 101 languages (Lu et al., 2024) with a total of 400M sentences. This work explores the impact of parallel data exclusively for low-resource performance and experiments with 1M–13M parallel sentences (50M–750M tokens) during pre-training, two orders of magnitude smaller than prior work.
@@
LLMs have been shown to perform poorly in low-resource MT (Robinson et al., 2023; Zhu et al., 2024b). In response, there have been efforts to leverage external resources in the MT pipeline, including multilingual lexicons (Lu et al., 2023), rule-based linguistic tools (Zhang et al., 2024b), word alignments (Mao and Yu, 2024) and even entire grammar books (Reid et al., 2024). However, such approaches create dependencies on resources and hinder extensibility across languages. Instead, we focus on optimal data utilisation strategies during CPT and SFT, prioritizing extensibility.
@@
There has been a body of work exploring multilingual instruction tuning that have touched on diversity and data scale, but most of it is limited to HRLs. MT examples were shown to improve cross-lingual generation (Ranaldi and Pucci, 2023; Zhu et al., 2023), while Chen et al. (2024) showed multilingual SFT on machine-translated Alpaca datasets matches or beats monolingual tuning. Kew et al. (2023) and Shaham et al. (2024) showed a small quantity of multilingual SFT data can improve cross-lingual generation capabilities in medium and high-resource languages, while Zhu et al. (2024a) showed only 32 examples in HRLs suffice to elicit MT capabilities from LLMs. In our work, we hypothesize that these findings have the common denominator of pre-training on HRLs at scale, and show that when moving to LRLs, these trends reverse and diversity is no longer beneficial.
@@ 3 Approach
We now describe our efforts to adapt the widely-used ALMA framework, originally designed for fine-tuning LLMs for HRL translation (Xu et al., 2024; Alves et al., 2024), for MT in LRLs.
@@ 3.1 Stage 1: Continued Pre-training (CPT)
In a scenario where monolingual data is scarce, it is crucial to investigate the most effective way to use parallel data—which, for our indigenous American languages, was found to surprisingly be more abundant than the former [^1] While monolingual online data is scarce, many translations of constitutions, articles etc. from Spanish do exist. We investigate 3 methods of mixing all available parallel and monolingual data:
@@
    All Mono: Here, we merge monolingual data with only the target side of all available bitext—essentially using it as extra monolingual data
@@
    Mono + parallel (concat): Here, we merge monolingual data with concatenated source-target pairs from parallel data. We prepend source and target language codes before concatenation and, following Guo et al. (2024), use a newline delimiter to separate them.
@@
    Mono + parallel (separate): To ablate the impact of concatenation, we provide the source and target sides of parallel data as separate sentences, and shuffle with monolingual data.
We depict our approach in Figure 1. In the first technique, the motivation is that it might be hard for the LLM to learn to model concatenated sequences, given that they were likely scarce in the original pre-training corpus, with the added challenge of pre-training on ‘new’ low-resource languages. On the other hand, if the model is able to adapt to concatenated sequences, it could make the LLM more adjusted to the task of translation. Finally, the third method verifies whether the results of the ‘concatenated’ model are due to the concatenation itself, rather than simply being exposed to additional tokens in the source language. We use this terminology for all experiments in this work.
@@
While these methods control how parallel data is incorporated in pre-training, we are also interested in adjusting the ratio of parallel to monolingual data in the corpus, particularly for the ‘concat’ method given pre-training on 100% parallel data can be suboptimal (Alves et al., 2024). To get a desired mixing ratio for ‘concat’, we include bitext only until it comprises a given percentage of the training corpus. Once this threshold is crossed, we use the ‘all mono’ method to include it as monolingual data instead. Next, we use temperature sampling (Arivazhagan et al., 2019) to control the language-wise distribution in our monolingual and parallel pre-training data, since these are quite heterogeneous and certain languages are extremely low-resourced. We sample monolingual and parallel data independently if using the ‘concat’ method, else we just mix them all together and shuffle at the instance level to create our final pre-training corpus.
@@ 3.2 Stage 2: Supervised Fine-Tuning (SFT)
Next, we fine-tune with LoRA on supervised instruction data and detail the tasks explored below. Note that we convert all instructions to the standard Alpaca format, and compute loss on the target tokens only (Taori et al., 2023).
@@
Apart from MT data in the LRLs, we also experiment with adding HRL MT data since it is more abundant and of higher quality, known to be important during SFT (Xu et al., 2024). To explore the impact of transfer learning, we work with HRL data that is in some way related to the source/target language, e.g. Spanish-English data for experiments on Spanish-X. Instructions are formatted in the same way as the LRL data.
@@
4 Experiments and Discussions
We fine-tune 2 separate sets of multilingual LLMs for 2 different language groups to facilitate evaluation on test sets from 2 different low-resource MT shared tasks: AmericasNLP 2024 (Ebrahimi et al., 2024) and the Indic track of WMT 2023 (Pal et al., 2023). The former involves 11 Indigenous Central & South American languages, while the latter focuses on 4 North-East (NE) Indian languages. The first group includes Aymara (aym), Bribri (bzd), Asháninka (cni), Chatino (ctp), Guarani (grn), Huichol (hch), Nahuatl (nhe), Otomi (ote), Quechua (quy), Shipibo-Konibo (shp) and Tarahumara (tar). The second consists of Khasi (kha), Meitei (mni), Mizo (lus) and Assamese (asm). Our motivation in choosing these languages was to experiment with LRLs containing both Latin (American) and non-Latin (Indic) scripts; that also had widely used, high-quality test sets. We use the former for our main experiments and replicate the most interesting baselines in the Indic languages.
@@ 4.1 Data
Table 3: Parallel data sentence counts for American languages, from source to each target language. HRL and LRL refer to the 3 high-resource languages (Aymara, Guarani, Quechua) and the other 8 low-resource ones respectively.
@@
We curate parallel data from various sources, for use in both CPT and SFT. Tables 3 and 4 show the sizes for the American and Indian languages respectively. Given that our primary exploration is for the American languages with limited spa-X data, we also sample eng-X and por-X from OPUS (Tiedemann, 2012). Note the heavily skewed language distribution, with the 3 HRLs constituting  80% of spa-X data and 96% of the overall data. The lesser skew for spa-X is due to the efforts of AmericasNLP to collect data for these pairs and the prevalence of Spanish in Latin American countries. A similar skew also exists for the Indic languages, with English-Khasi being the least-resourced pair. We list sources for all curated data, along with cleaning steps, in Appendix A.